OpenStudy (anonymous):

b. The water line is given by the equation 12 3 2 y = - x - What is the slope and y-intercept of the line in Part b? How do you know?

OpenStudy (saifoo.khan):

please rewrite.q

OpenStudy (angela210793):

can u pls write the eq once again :)

OpenStudy (anonymous):

y=-2/3x - 12

OpenStudy (angela210793):

y=mx+t (m=slope) slope is -2/3 and u intercept is -12

OpenStudy (katrinakaif):

slope -2/3 y intercept -12 y=mx + b m = slope b = yintercept